if one sphere has a mass of 2 kg in electric charge power 4 microcoulomb while another sphere 3 metres away is a mass of 1 kg in electric charge of 5 microcoulomb what is the acceleration of the more massive sphere
1
Expert's answer
2020-08-13T12:23:37-0400

The expression for the electrostatic force between two charges is

"F=\frac{kq_1q_2}{r^2}=\frac{(9*10^9)(4*10^{-6})(5*10^{-6})}{(3)^2}"

"=0.02 N"

So the acceleration of the more massive sphere is

"a=\frac{F}{m}=\frac{0.02 N}{2 kg}=0.01 m/s^2"